If 65g of Aluminum reacts with excess Chlorine, how many grams of Aluminum Chloride will form?


1
Expert's answer
2021-10-29T02:19:19-0400

2Al + 3Cl2 → 2AlCl3

MM(Al) = 26.98 g/mol

n(Al) "= \frac{65}{26.98} = 2.41 \;mol"

According to the reaction:

n(AlCl3) =n(Al) = 2.41 mol

MM(AlCl3) = 133.34 g/mol

m(AlCl3) "= 2.41 \times 133.34 = 321.35 \;g"